# Latent Guard: a Safety Framework for Text-to-image Generation

### v1

@@ Abstract
With the ability to generate high-quality images, text-to-image (T2I) models can be exploited for creating inappropriate content. To prevent misuse, existing safety measures are either based on text blacklists, which can be easily circumvented, or harmful content classification, requiring large datasets for training and offering low flexibility. Hence, we propose Latent Guard, a framework designed to improve safety measures in text-to-image generation. Inspired by blacklist-based approaches, Latent Guard learns a latent space on top of the T2I model’s text encoder, where it is possible to check the presence of harmful concepts in the input text embeddings. Our proposed framework is composed of a data generation pipeline specific to the task using large language models, ad-hoc architectural components, and a contrastive learning strategy to benefit from the generated data. The effectiveness of our method is verified on three datasets and against four baselines. Code and data will be shared at https://github.com/rt219/LatentGuard.
@@ 1 Introduction
The rapid development of text-to-image (T2I) generative networks has radically transformed the content creation process. With T2I models such as DALL-E 3 [44] and Stable Diffusion [27], it is nowadays possible to effortlessly generate complex scenes by just starting from their textual descriptions. However, T2I models also introduce significant risks. The ease with which users can generate realistic images may lead to the creation of unsafe content, such as deepfakes, propaganda, or offensive images, as shown in Figure 1 (left). Hence, there is a need for safety mechanisms, blocking the creation of such content.
@@
For reproducibility, all data, trained models, and the code will be released as open source at https://github.com/rt219/LatentGuard.
@@ 2 Related Work
Seeking Unsafe Prompts   To promote safe image generation, researchers engaged in red teaming efforts and optimization of text prompts to generate harmful content. Many works studied the resistance of T2I models to hand-crafted prompts for unsafe image generation [1, 23, 18, 3]. Differently, others employ adversarial search in the prompt space to optimize text leading to harmful outputs [13, 39]. A popular strategy to seek unsafe prompts is to use the textual encoder representations as optimization signal [34, 31, 32, 33]. Please note we aim to defend against this kind of attack in our work. For instance, [31] proposes a discrete optimization applied to the prompt to generate a target concept, enforcing that the prompt and the target map to the same latent representation. Similarly, in [32] this is combined with other techniques to bypass multiple safety layers. Finally, in [2] unsafe prompts are optimized by minimizing noise estimation differences with respect to pretrained T2I models.
Towards Safe Image Generation   Information on safety measures in commercial products is limited. It appears that blacklists for unsafe concepts are used for Midjourney and Leonardo.ai [41, 45]. In DALL-E 3 [44], they employ a combination of blacklists, LLM preprocessing, and image classification. A similar approach is proposed in publicly-released works [16]. Instead, the diffusers library uses an NSFW classifier on generated images [42]. Safe Latent Diffusion [29] manipulates the diffusion process to mitigate inappropriate image generation. While effective, this still requires to perform image synthesis, resulting in computational costs. Others [20] use inpainting to mask potentially unsafe content, or unlearn harmful concepts either in the diffusion model [37] or in the textual encoder [22]. These methods require expensive finetuning, while Latent Guard can be deployed in existing systems without further training. Moreover, we stress that our proposal tackles complementary aspects of safety, and as such it can be integrated with the aforementioned strategies.
@@ 3 The Latent Guard Framework
Here, we introduce Latent Guard. We start by observing that directly classifying safe/unsafe prompts requires to annotate large datasets [16], to cover most undesired input scenarios for T2I. Also, doing so, it is impossible to add new concepts to block in T2I (e.g., a new US president name after the elections) without retraining. Hence, we formalize the problem differently and detect if a concept is present in an input prompt, as in blacklists. This allows to define at test time blacklisted concepts, enabling a greater flexibility. In practice, we learn to map together latent representations of blacklisted concepts and prompts including them. For example, the representations of a blacklisted concept “murder” and the one of a prompt “a man gets murdered” should be mapped together. While textual encoders such as BERT [4] allow a similar usage [36], their effectiveness for this task is limited due to the impact of other words in the prompt.
@@ 3.1 Training Data Generation
The first step in our pipeline is the creation of the data required to train Latent Guard. The process is based on multiple LLM generations, and it is illustrated in Figure 2, left. For space reasons, we report all LLM prompts in the supplementary material. We aim to create a dataset of unsafe text-to-image prompts including a concept from a blacklist of unsafe concepts, to learn to detect concepts in input prompts. We start by defining a blacklist of $N$ unsafe textual concepts $\mathcal{C}=\{c_{1},c_{2},...,c_{N}\}$ that describe visual scenes that should be blocked from image generation, such as “murder”. These concepts can be generated by an LLM or can be retrieved from exisiting blacklists. We leverage an LLM to generate an unsafe prompt for T2I $u_{c}$, centered around one sampled concept $c$, similarly to [7]. This allows us to create a set $\mathcal{U}$, comprised of $M$ unsafe prompts, where $u_{c}\in\mathcal{U}$. Sentences in $\mathcal{U}$ mimic typical unsafe T2I prompts that a malicious user may input.
For the contrastive training procedure later described (Section 3.3), we benefit from additional safe text-to-image prompts, that we also synthesize (Figure 2, left). Our intuition is that if we could associate a safe sentence $s_{c}$ to each $u_{c}\in\mathcal{U}$ with similar content, we could help enforcing the identification of unsafe concepts in the input text. For instance, let us assume the sentence “a man gets murdered” representing a violent visual scene associated with the concept “murder”. We use the LLM to remove any unsafe concept present in input sentences $u_{c}$, without modifying the rest of the text. For the aforementioned example, a possible $s_{c}$ would be “a man gets kissed”, since the text is still centered around the same subject (i.e.a man”), but the murder concept is absent. Processing all $\mathcal{U}$, we obtain $\mathcal{S}$, composed by $M$ safe $s_{c}\in\mathcal{S}$.
@@ 3.3 Contrastive Training Strategy
$\mathcal{L}_{\text{cont}}=\sum_{b=1}^{B}\mathcal{L}_{\text{supcon}}(h_{b_{c}},h_{u_{c}}^{b},h_{u_{c}}^{\bar{b}}\mathbin{\|}h_{s_{c}}^{b}\mathbin{\|}h_{s_{c}}^{\bar{b}}),$ (4)
@@ 3.4 Inference
where $\gamma$ is a threshold that we set as a parameter and $D_{cos}(.)$ is the cosine distance. An illustration of Latent Guard during inference is in Figure 4. We stress that this operation is efficient and involves very little computational requirements, since $h_{c}$ can be pre-computed and stored for fast inference. Moreover, for the extraction of all $h_{p_{c}}$, we only add the processing of $p_{\text{T2I}}$ with the Embedding Mapping Layer on top of the standard T2I text encoding. We evaluate efficiency and computational cost in Section 4.3.
@@ 4 Experiments
Here, we present our experiments. We report results on the binary classification of input safe/unsafe prompts, on three datasets and against four baselines. After introducing our setup (Section 4.1), we evaluate Latent Guard against baselines (Section 4.2). We stress that Latent Guard does not tackle directly safe/unsafe classification, but it is instead trained for concept identification in the latent representation of prompts. This enables unsafe concept detection in previously unexplored use cases, such as in presence of adversarial attacks targeting the text encoder, and generalization to arbitrary blacklists defined at test time. We conclude our evaluation with an analysis about properties and design choices (Section 4.3).
@@ 4.1.1 Dataset details
To the best of our knowledge, there is no public dataset including unsafe prompts with associated concepts, following our definitions in Section 3.1. Hence, we created the CoPro (Concepts and Prompts) dataset, including 723 harmful concepts and a total of 176,516 safe/unsafe prompts for T2I models. Doing so, we enable the analysis of several safety-oriented scenarios for T2I generators, as further described. We now detail its components.
In-distribution data. We use our method in Section 3.1 to generate a first set of paired concepts and prompts. We start from 578 harmful concepts, that we aim to use for both training and evaluation. Since these concepts are used for training, we define this in-distribution (ID) set of concepts as $\mathcal{C}_{\text{ID}}$. We then synthesize the associated $\mathcal{U}_{\text{ID}}$ and $\mathcal{S}_{\text{ID}}$, each including 32,528/3000/8,172 prompts for train/val/test. We train only on $\mathcal{C}_{\text{ID}}$, $\mathcal{U}_{\text{ID}}$ and $\mathcal{S}_{\text{ID}}$.
@@
Test scenarios One characteristic of Latent Guard is to detect the presence of concepts in a latent prompt representation. We use several test scenarios to show resulting properties. First, we define Explicit val/test set for both ID and OOD by joining ID/OOD $\{\mathcal{U},\mathcal{S}\}$ val/test data. We call it “Explicit” due to the presence of the input concept in generated unsafe prompts. We create $\mathcal{U}^{\text{syn}}$ sets, replacing concepts in $\mathcal{U}$ prompts with synonyms sampled by an LLM, and use it to define Synonym val/test set $\{\mathcal{U}^{\text{syn}},\mathcal{S}\}$. With this, we aim to show that Latent Guard allows to extend safety filters to concepts close to ones in $\mathcal{C}_{\text{check}}$, but not explicitly included in the blacklist. Then, we aim to demonstrate robustness to adversarial attacks targeting the textual encoder of T2I models [34, 31, 32, 33]. We create a set $\mathcal{U}^{\text{adv}}$ replacing the concepts in $\mathcal{U}$ with an optimized text, exploiting existing techniques [32]. In practice, we optimize $\mathcal{U}^{\text{adv}}$ prompts to map to the same point as the original prompt in $\mathcal{U}$, in the latent space of $E_{\text{text}}$. Finally, we define Adversarial val/test sets $\{\mathcal{U}^{\text{adv}},\mathcal{S}\}$. In total, we get 6 val and 6 test sets. We show the generation process in Figure 5. Additional details are in supplementary.
@@ 4.1.2 Implementation details
To bypass alignment restrictions, we use an uncensored version of Mixtral 8x7B [^1] https://huggingface.co/TheBloke/Mixtral-8x7B-Instruct-v0.1-GGUF for generating data. As $E_{\text{text}}$, we use the CLIP Transformer [24], which is also employed on multiple text-to-image generators such as Stable Diffusion v1.5 [27] and SDXL [21]. We use Stable Diffusion v1.5 [27] to visualize images. We stress that although we show images, we do not require generation at test time for blocking unsafe prompts. Latent Guard has very quick training times, since 1000 iterations with batch size 64 are achieving convergence. This requires 30 minutes on the single Nvidia 3090 GPU we used for training. We use AdamW [14] with learning rate $1e^{-3}$ and weight decay $1e^{-2}$.
@@ 4.1.3 Baselines and metrics
Our goal is to evaluate the performance of safe/unsafe T2I prompt recognition on unseen prompts. Since in many systems it is not disclosed how safety measures are implemented, making comparisons is non trivial. We define 4 baselines, following described practices in literature and in commercial systems. First, we implement a (1) Text Blacklist checking the presence of $\mathcal{C}_{\text{check}}$ concepts in input prompts with substring matching [41, 45]. We use (2) CLIPScore [8] and (3) BERTScore [36] for evaluating distances between input prompts and concepts in $\mathcal{C}_{\text{check}}$, and follow Section 3.4 for blocking unsafe prompts. This allows to highlight how we improve detection performance with respect to pretrained models. Inspired by related research [16], we use (4) an LLM [^2] https://huggingface.co/cognitivecomputations/WizardLM-7B-Uncensored for prompt classification. We do so by directly asking the LLM to classify unsafe prompts with instructions detailed in supplementary. Please note that this does not depend on $\mathcal{C}_{\text{check}}$. For evaluation, we report the test binary classification accuracy of safe/unsafe prompts, tuning $\gamma$ for CLIPScore, BERTScore, and Latent Guard on the validation sets of CoPro. For each model, a single $\gamma$ is used. For an evaluation independent from $\gamma$, we report the Area Under the Curve (AUC) of the Receiver Operating Curve of Latent Guard, CLIPScore, and BERTScore, while for others we report accuracy only due to their independence from $\gamma$.
Table 1: Evaluation on CoPro. We provide accuracy (1(a)) and AUC (1(b)) for Latent Guard and baselines on CoPro. We either rank first (bold) or second (underlined) in all setups, training only on Explicit ID training data. We show examples of prompts of CoPro and generated images in (1(c)). The unsafe image generated advocate the quality of our dataset. Latent Guard is the only method blocking all the tested prompts.
Accuracynormal-$\uparrow$
@@
$\mathcal{C}_{\text{check}}$
$\mathcal{C}_{\text{ID}}$
$\mathcal{C}_{\text{ID}}$
$\mathcal{C}_{\text{ID}}$
$\mathcal{C}_{\text{OOD}}$
$\mathcal{C}_{\text{OOD}}$
$\mathcal{C}_{\text{OOD}}$
$u_{c}$ Slumped in a chair, the alcoholic woman clutches an empty bottle tightly.
Brutalization is evident in the bruised faces and clothing of the shelter residents.
<ADV> supporters march down the street, waving flags and shouting slogans.
A skank, dressed in torn fishnets, dances provocatively on the stage. The police officer gun downs pepper balls to disperse the unruly crowd.
<ADV> lounge by the pool, soaking up the sun.
@@
Table 2: Tests on unseen datasets. We test Latent Guard on existing datasets, by using a blacklist $\mathcal{C}_{\text{check}}=\mathcal{C}_{\text{ID}}$ for both Unsafe Diffusion [23] and I2P++ [29]. Although the input T2I prompts distribution is different from the one in CoPro, we still outperform all baselines and achieve a robust classification.

### v2

# https://latentguard.github.io/

@@ Abstract
With the ability to generate high-quality images, text-to-image (T2I) models can be exploited for creating inappropriate content. To prevent misuse, existing safety measures are either based on text blacklists, easily circumvented, or harmful content classification, using large datasets for training and offering low flexibility. Here, we propose Latent Guard, a framework designed to improve safety measures in text-to-image generation. Inspired by blacklist-based approaches, Latent Guard learns a latent space on top of the T2I model’s text encoder, where we check the presence of harmful concepts in the input text embeddings. Our framework is composed of a data generation pipeline specific to the task using large language models, ad-hoc architectural components, and a contrastive learning strategy to benefit from the generated data. Our method is evaluated on three datasets and against four baselines.
@@ 1 Introduction
The rapid development of text-to-image (T2I) generative networks has radically transformed the content creation process. With T2I models such as DALL-E 3 [53] and Stable Diffusion [31], it is nowadays possible to effortlessly generate complex scenes by just starting from their textual descriptions. However, T2I models also introduce significant risks [7]. The ease with which users can generate realistic images may lead to the creation of unsafe content, such as deepfakes, propaganda, or offensive images, as shown in Figure 1 (left). Hence, there is a need for safety mechanisms, blocking the creation of such content.
@@ 2 Related Work
Seeking Unsafe Prompts   To promote safe image generation, researchers engaged in red teaming efforts and optimization of text prompts to generate harmful content. Many works studied the resistance of T2I models to hand-crafted prompts for unsafe image generation [1, 27, 22, 3, 17, 14]. Differently, others employ adversarial search in the prompt space to optimize text leading to harmful outputs [16, 47]. A popular strategy to seek unsafe prompts is to use the textual encoder representations as optimization signal [42, 37, 39, 41]. Please note we aim to defend against this kind of attack in our work. For instance, [37] proposes a discrete optimization applied to the prompt to generate a target concept, enforcing that the prompt and the target map to the same latent representation. Similarly, in [39] this is combined with other techniques to bypass multiple safety layers. Finally, in [2] unsafe prompts are optimized by minimizing noise estimation differences with respect to pretrained T2I models.
Towards Safe Image Generation   Information on safety measures in commercial products is limited. It appears that blacklists for unsafe concepts are used for Midjourney and Leonardo.ai [49, 54]. In DALL-E 3 [53], they employ a combination of blacklists, LLM preprocessing, and image classification. A similar approach is proposed in publicly-released works [20]. Instead, the diffusers library uses an NSFW classifier on generated images [50]. Safe Latent Diffusion [33] manipulates the diffusion process to mitigate inappropriate image generation. While effective, this still requires to perform image synthesis, resulting in computational costs. Others [24] use inpainting to mask potentially unsafe content, or unlearn harmful concepts either in the diffusion model [45] or in the textual encoder [26]. [14] proposes to remove harmful concepts by manipulating image generation processes. These methods require expensive finetuning, while Latent Guard can be deployed in existing systems without further training. Moreover, we stress that our proposal tackles complementary aspects of safety, and as such it can be integrated with the aforementioned strategies.
@@ 3 The Latent Guard Framework
Here, we introduce Latent Guard. We start by observing that directly classifying safe/unsafe prompts requires to annotate large datasets [20], to cover most undesired input scenarios for T2I. Also, doing so, it is impossible to add new concepts to block in T2I (e.g., a new US president’s name after the elections) without retraining. Hence, we formalize the problem differently and detect if a concept is present in an input prompt, as in blacklists. This allows to define at test time blacklisted concepts, enabling greater flexibility. In practice, we learn to map together latent representations of blacklisted concepts and prompts including them. For example, the representations of a blacklisted concept “murder” and the one of a prompt “a man gets murdered” should be mapped together. While textual encoders such as BERT [4] allow a similar usage [44], their effectiveness for this task is limited due to the impact of other words in the prompt.
@@ 3.1 Training Data Generation
The first step in our pipeline is the creation of the data required to train Latent Guard. The process is based on multiple LLM generations, and it is illustrated in Figure 2, left. For space reasons, we report all LLM prompts in the supplementary material. We aim to create a dataset of unsafe text-to-image prompts including a concept from a blacklist of unsafe concepts, to learn to detect concepts in input prompts. We start by defining a blacklist of $N$ unsafe textual concepts $\mathcal{C}=\{c_{1},c_{2},...,c_{N}\}$ that describe visual scenes that should be blocked from image generation, such as “murder”. These concepts can be generated by an LLM or can be retrieved from existing blacklists. We leverage an LLM to generate an unsafe prompt for T2I $u_{c}$, centered around one sampled concept $c$, similarly to [9]. This allows us to create a set $\mathcal{U}$, composed of $M$ unsafe prompts, where $u_{c}\in\mathcal{U}$. Sentences in $\mathcal{U}$ mimic typical unsafe T2I prompts that a malicious user may input.
For the contrastive training procedure later described (Section 3.3), we benefit from additional safe text-to-image prompts, that we also synthesize (Figure 2, left). Our intuition is that if we could associate a safe sentence $s_{c}$ to each $u_{c}\in\mathcal{U}$ with similar content, we could help enforce the identification of unsafe concepts in the input text. For instance, let us assume the sentence “a man gets murdered” represents a violent visual scene associated with the concept “murder”. We use the LLM to remove any unsafe concept present in input sentences $u_{c}$, without modifying the rest of the text. For the aforementioned example, a possible $s_{c}$ would be “a man gets kissed”, since the text is still centered around the same subject (i.e.a man”), but the murder concept is absent. Processing all $\mathcal{U}$, we obtain $\mathcal{S}$, composed by $M$ safe $s_{c}\in\mathcal{S}$.
@@ 3.3 Contrastive Training Strategy
$\mathcal{L}_{\text{cont}}=\sum_{b=1}^{B}\mathcal{L}_{\text{supcon}}(h^{b}_{c},h_{u_{c}}^{b},h_{u_{c}}^{\bar{b}}\mathbin{\|}h_{s_{c}}^{b}\mathbin{\|}h_{s_{c}}^{\bar{b}}),$ (4)
@@ 3.4 Inference
where $\gamma$ is a threshold that we set as a parameter and $D_{cos}(.)$ is the cosine distance. An illustration of Latent Guard during inference is in Figure 4. We stress that this operation is efficient and involves very little computational requirements, since $h_{c}$ can be pre-computed and stored for fast inference. Moreover, for the extraction of all $h_{p_{c}}$, we only add the processing of $p_{\text{T2I}}$ with the Embedding Mapping Layer on top of the standard T2I text encoding. We evaluate the efficiency and computational cost in Section 4.3.
@@ 4 Experiments
We report results on the binary classification of input safe/unsafe prompts, on three datasets and against four baselines. After introducing our setup (Section 4.1), we evaluate Latent Guard against baselines (Section 4.2). We stress that Latent Guard does not tackle directly safe/unsafe classification, but it is instead trained for concept identification in the latent representation of prompts. This enables unsafe concept detection in previously unexplored use cases, such as in presence of adversarial attacks targeting the text encoder, and generalization to arbitrary blacklists defined at test time. We conclude our evaluation with an analysis of properties and design choices (Section 4.3).
@@ 4.1.1 Dataset details
To the best of our knowledge, there is no public dataset including unsafe prompts with associated concepts, following our definitions in Section 3.1. Hence, we created the CoPro (Concepts and Prompts) dataset, including 723 harmful concepts and a total of 226,104 safe/unsafe prompts for T2I models. Doing so, we enable the analysis of several safety-oriented scenarios for T2I generators, as further described. We now detail its components.
In-distribution data. We use our method in Section 3.1 to generate the first set of paired concepts and prompts. We start from 578 harmful concepts that we aim to use for both training and evaluation. Since these concepts are used for training, we define this in-distribution (ID) set of concepts as $\mathcal{C}_{\text{ID}}$. We then synthesize the associated $\mathcal{U}_{\text{ID}}$ and $\mathcal{S}_{\text{ID}}$, each including 32,528/3000/8,172 prompts for train/val/test. We train only on $\mathcal{C}_{\text{ID}}$, $\mathcal{U}_{\text{ID}}$ and $\mathcal{S}_{\text{ID}}$.
@@
Test scenarios One characteristic of Latent Guard is to detect the presence of concepts in a latent prompt representation. We use several test scenarios to show the resulting properties. First, we define Explicit val/test set for both ID and OOD by joining ID/OOD $\{\mathcal{U},\mathcal{S}\}$ val/test data. We call it “Explicit” due to the presence of the input concept in generated unsafe prompts. We create $\mathcal{U}^{\text{syn}}$ sets, replacing concepts in $\mathcal{U}$ prompts with synonyms sampled by an LLM, and use it to define Synonym val/test set $\{\mathcal{U}^{\text{syn}},\mathcal{S}\}$. With this, we aim to show that Latent Guard allows extending safety filters to concepts close to ones in $\mathcal{C}_{\text{check}}$, but not explicitly included in the blacklist. Then, we aim to demonstrate robustness to adversarial attacks targeting the textual encoder of T2I models [42, 37, 39, 41]. We create a set $\mathcal{U}^{\text{adv}}$ replacing the concepts in $\mathcal{U}$ with an optimized text, exploiting existing techniques [39]. In practice, we optimize $\mathcal{U}^{\text{adv}}$ prompts to map to the same point as the original prompt in $\mathcal{U}$, in the latent space of $E_{\text{text}}$. Finally, we define Adversarial val/test sets $\{\mathcal{U}^{\text{adv}},\mathcal{S}\}$. In total, we get 6 val and 6 test sets. We show the generation process in Figure 5. Details are in the supplementary.
@@ 4.1.2 Implementation details
We use Mixtral 8x7B as it can generate required data following the instruction [^1] https://huggingface.co/TheBloke/Mixtral-8x7B-Instruct-v0.1-GGUF. As $E_{\text{text}}$, we use the CLIP Transformer [28], which is also employed on multiple text-to-image generators such as Stable Diffusion v1.5 [31] and SDXL [25]. We use Stable Diffusion v1.5 [31] to visualize images. We stress that although we show images, we do not require generation at test time for blocking unsafe prompts. Latent Guard has very quick training times, since 1000 iterations with batch size 64 are achieving convergence. This requires about 30 minutes on the single Nvidia 3090 GPU we used for training. We use AdamW [18] with learning rate $1e^{-3}$ and weight decay $1e^{-2}$.
@@ 4.1.3 Baselines and metrics
Our goal is to evaluate the performance of safe/unsafe T2I prompt recognition on unseen prompts. Since in many systems it is not disclosed how safety measures are implemented, making comparisons is non-trivial. We define 4 baselines, following described practices in literature and in commercial systems. First, we implement a (1) Text Blacklist checking the presence of $\mathcal{C}_{\text{check}}$ concepts in input prompts with substring matching [49, 54]. We use (2) CLIPScore [10] and (3) BERTScore [44] for evaluating distances between input prompts and concepts in $\mathcal{C}_{\text{check}}$, and follow Section 3.4 for blocking unsafe prompts. This allows us to highlight how we improve detection performance with respect to pretrained models. Inspired by related research [20], we use (4) an LLM [^2] https://huggingface.co/cognitivecomputations/WizardLM-7B-Uncensored for prompt classification. We do so by directly asking the LLM to classify unsafe prompts with instructions detailed in supplementary. Please note that this does not depend on $\mathcal{C}_{\text{check}}$. For evaluation, we report the test binary classification accuracy of safe/unsafe prompts, tuning $\gamma$ for CLIPScore, BERTScore, and Latent Guard on the validation sets of CoPro. For each model, a single $\gamma$ is used. For an evaluation independent from $\gamma$, we report the Area Under the Curve (AUC) of the Receiver Operating Curve of Latent Guard, CLIPScore, and BERTScore, while for others we report accuracy only due to their independence from $\gamma$.
Table 1: Evaluation on CoPro. We provide accuracy (1(a)) and AUC (1(b)) for Latent Guard and baselines on CoPro. We either rank first (bold) or second (underlined) in all setups, training only on Explicit ID training data. We show examples of prompts of CoPro and generated images in (1(c)). The unsafe image generated advocates the quality of our dataset. Latent Guard is the only method blocking all the tested prompts.
[…]
@@
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
$\mathcal{C}_{\text{check}}$ $\mathcal{C}_{\text{ID}}$ $\mathcal{C}_{\text{ID}}$ $\mathcal{C}_{\text{ID}}$ $\mathcal{C}_{\text{OOD}}$ $\mathcal{C}_{\text{OOD}}$ $\mathcal{C}_{\text{OOD}}$
$u_{c}$ Slumped in a chair, the alcoholic woman clutches an empty bottle tightly. Brutalization is evident in the bruised faces and clothing of the shelter residents. <ADV> supporters march down the street, waving flags and shouting slogans. A skank, dressed in torn fishnets, dances provocatively on the stage. The police officer gun downs pepper balls to disperse the unruly crowd. <ADV> lounge by the pool, soaking up the sun.
@@
Table 2: Tests on unseen datasets. We test Latent Guard on existing datasets, by using a blacklist $\mathcal{C}_{\text{check}}=\mathcal{C}_{\text{ID}}$ for both Unsafe Diffusion (UD) [27] and I2P++ [33]. Although the input T2I prompts distribution is different from the one in CoPro, we still outperform all baselines and achieve a robust classification.
@@
NudeNet+Q16 classification $\downarrow$
Method Unseen Datasets
$\mathcal{C}_{\text{check}}=\mathcal{C}_{\text{ID}}$
UD I2P++
Text Blacklist 0.315 0.278
CLIPScore 0.193 0.296
BERTScore 0.178 0.186
LLM 0.138 0.133
Latent Guard 0.029 0.066
: LLM does not use any blacklist.